OpenStudy (anonymous):

In an egg carton there are 12 eggs, of which 9 are hard-boiled and 3 are raw. Six of the eggs are chosen at random to take to a picnic (yes, the draws are made without replacement). Find the chance that at least one of the chosen eggs is raw.

OpenStudy (anonymous):

its 10/11 it has been explained in the previous attempt

OpenStudy (anonymous):

total no. of eggs =12 A=eggs are half boiled B=eggs are raw P(B)=3/12=1/4 P(not B) =P'=1-1/4=3/4 Probability no raw egg=3/4*3/4*3/4*3/4*3/4*3/4=729/4096 Probability at least one raw egg=1-729/4096=3367/4096